HUBBLE SPACE TELESCOPE DAILY REPORT #2308 PERIOD COVERED: 0000Z (UTC) 02/12/99 - 0000Z (UTC) 02/15/99 Daily Status Report as of 046/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Six Sets of STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Eight Sets of WF/PC-2 7713 (WF/P-2 Cycle 7 Supplemental Darks pt3/3) The WF/PC-2 was used to make supplemental dark calibration images. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no anomalies were noted. 1.3 Completed STIS/CCD/MA1 7263 (Physical Processes in Stellar Atmospheres: Comparative Analysis of the Sun and Alpha Cen A) The Space Telescope Imaging Spectrograph (CCD and MA1) were used to study the physical processes in the stellar atmosphere of Alpha-Cen (a) (a.k.a., HD128620). Using this atmosphere along with observed velocity fields, we will reassess competing theories for various phenomena observed in solar-type stars. A wavelength calibration image was also taken. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed Seven Sets of WF/PC-2 8090 (POMS Test Proposal: WF-II Parallel Archive Proposal Continuation) The WF/PC-2 was used in the parallel mode to make some low galactic latitude archive observations. The program will be used to take parallel images of random areas of the sky. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed Two Sets of STIS/CCD 7475 (The Chemical Content of the Universe at Z< 1.6) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the background quasar Q0454+0356 and Field-Q093. Damped Lyman-Alpha absorption systems provide a means of directly probing the chemical content of the deep universe. A wavelength calibration image was also taken. There was a Guide Star Acquisition which resulted in a Fine Lock Backup at 045/1423Z due to a FGS #1 SSLE. The subsequent reacquisiton (1,3,1) resulted in fine lock backup, Fine lock was achieved using the same primary FGS #1 (1,0,1). HSTART 6997 was written. No other problems were reported. 1.6 Completed WF/PC-2 8061 (POMS Test Proposal: WF-II Backup Parallel Archive Proposal) This is a POMS test proposal designed to simulate scientific plans. The proposal was executed as planned, and no anomalies were reported. 1.7 Completed STIS/CCD 7297 (Understanding the Red Rectangle and its Central Binary HD 44179) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the Red Rectangle and its central binary star HD44179. The Red Rectangle is one of the nearest and best known of the proto-planetary nebulae. A wavelength calibration observation was also taken. The observations were completed as planned, and no problems were reported. 1.8 Completed WF/PC-2 7386 (A Snapshot Survey of Nova Shells) The WF/PC-2 was used to make observational images of the super nova shell Nova-Lup-1 in order to detect and measure its expanding gaseous shell. These observations will trace the distribution of hot and cool gas in the nova shells, and show their degree of inhomogeneity on both large and small scales. The observations were completed as scheduled, and no problems were noted. 1.9 Completed Twelve Sets of STIS/CCD 8091 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make high galactic latitude, non-scripted, parallel observations. There was a Guide Star Acquisition which resulted in a Fine Lock Backup at 045/1423Z due to a FGS #1 SSLE. The subsequent reacquisiton (1,3,1) resulted in fine lock backup, Fine lock was achieved using the same primary FGS #1 (1,0,1). HSTART 6997 was written. no other anomalies were noted. 1.10 Completed WF/PC-2 7447 (The Cosmological Evolution of Quasar Host Galaxies) The WF/PC-2 was used to make observations of the host galaxy which surrounds quasar 1239-041 in order to study it cosmological evolution. The observations were executed as scheduled, and no anomalies were reported. 1.11 Completed Three Sets of WF/PC-2 6353 (A Snapshot Emission Line Imaging Survey of Very Low Excitation Planetary Nebulas) The WF/PC-2 was used to make snapshot observations as part to an emission line survey of the very low excitation planetary nebula PK285-02D1, PK002-03D3 and PK326-06D1 (Class PNe). This target was selected based on the O III/H alpha line flux ratio. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed FOC/96/WF/PC-2 7274 (Secular Changes in the Jet of M-87) The Faint Object Camera (f/96) and the WF/PC-2 were both used to make observations of the secular changes in the jet coming off of the active galactic nuclei of galaxy M-87. Past observations with the FOC have shown changes as small as 0.03", and provide a unique opportunity to study proper motions and structural changes of the jet. An internal flat calibration image was also taken. The observations were completed as scheduled, and no anomalies were reported. 1.13 Completed WF/PC-2 6357 (The Nuclear Morphology of Elliptical Galaxies) The WF/PC-2 was used to make images of the early type, large, elliptical galaxy ESO376-9 (taken from a list of 150 galaxies) in order to study its nuclear morphology. This survey will complement Illingworth's snapshot study of later Hubble types. There was a guide star acquisition failure (3,0,3) at 044/0122Z. Neither the Scan Step or SRLE flags were set. HSTAR 6992 was written. No other anomalies were reported. 1.14 Completed Two Sets of STIS/MA2 7748 (The Origins of Boron) The Space Telescope Imaging Spectrograph (MA2) was used to make observations of the stars HD294297 and HD294297 which is part of the Orion OB1 class association in order to study the origins of Boron. The youngest stars show higher oxygen abundances than the oldest stars. A wavelength calibration image was also taken. The observations were completed as planned, and no problems were reported. 1.15 Completed STIS/CCD 7436 (A Complete Sample of Hot Post-AGB Stars in Globular Clusters) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of the globular cluster NGC-1904. Recent observations with the Ultraviolet Imaging Telescope have provided a complete sample of the hot post asymptotic giant branch {post-AGB} population in fourteen globular clusters. A wavelength calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 1.16 Completed STIS/CCD/WF/PC-2 7367 (Multi-Wavelength Imaging of Circumstellar Disks in the Orion Nebula) The Space Telescope Imaging Spectrograph (CCD) and the WF/PC-2 were used to make multi-wavelength imaging of circumstellar disks (NGC-1976) seen as silhouettes against the bright emission-line background in the Orion Nebula. With these data one should be able to examine the radial structure of the disks in great detail, and over a wide spatial extent. The observations were completed as planned, and no problems were reported. 1.17 Completed Two Sets of WF/PC-2 7387 (T-Tauri Star Snapshot Survey) The WF/PC-2 was used to make observations of the stars SZ-68 and RU-Lup as part of a T-Tauri class star survey. The observations were executed as scheduled, and no anomalies were noted. 1.18 Completed Two Sets of STIS/CCD 8084 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation) The Space Telescope Imaging Spectrograph (CCD) was used to make non-scripted parallel high galactic latitude observations as part of POMS test proposal. The observations were executed as scheduled, and no anomalies were noted. 1.19 Completed WF/PC-2 7377 (Dwarf Elliptical Galaxy Snapshot Survey II) The WF/PC-2 was used to make observations of the dwarf elliptical galaxy VCC1448 as part of Snapshot Survey II. This survey will provide important information on the globular cluster systems of such galaxies and the properties of their cores and nuclei. The observations were completed as planned, and no anomalies were reported. 1.20 Completed Three Sets of WF/PC-2 7505 (The Hubble Diagram for Distant Supernovae Measuring Cosmic Deceleration and Global Curvature) The WF/PC-2 was used to make observations of Super Nova SN1998 C1, A1 and A2 in order to produce a precise Hubble diagram from measuring cosmic deceleration and global curvature. HST can make an essential contribution by providing photometry with substantially higher precision and better temporal coverage than can be obtained from the ground. Before the start of the observations the Gyro #1 Saturation Safing Flag was out of limits (OOL) for one sample (12 Seconds) at the end of a slew at 045/1033Z. HSTAR 6996 was written. No other anomalies were reported. 1.21 Completed WF/PC-2 6458 (Imaging of QSO Host Galaxies at Z >2) The WF/PC-2 was used to make observations of the QSO/host galaxy 1338+277. QSOs of redshift larger than 2 have been resolved with ~arcsecond resolution on 4m-class ground-based telescopes. Both line and continuum flux have been detected over a few arcseconds, but the nature of the extended emission, and its relation to the nuclear energy source is unclear. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 24 Successful : 23 There was a guide star acquisition failure (3,0,3) at 044/0122Z. Neither the Scan Step or SRLE flags were set. HSTAR 6992 was set. See Proposal 6357 (Section 1.13 of this report). Scheduled Reacquisitions : 27 Successful : 27 2.2 FHST Updates: Scheduled : 51 Successful : 51 2.3 Operations Notes: The NICMOS FPA Temerature Sensor Dectector #2 (NFPA2TMP) was observed to have a value > 300 Degrees (K) at 045/2040Z. Per Ops Note #552, OR 14292 was executed to re-enable the EPA temperature sensor at 045/2307Z. The MOSES SISE was notified. HSTAR 6998 was written. There were twenty five missing minor frames of data at 046/0516Z (TTR 20588). There was a data inversion (TTR 17864) at 046/0522Z, 0526Z, and 0543Z (generic TTR). A SIP was executed at 043/2032Z (OR 14312). A NSSC-1 Status Buffer was dumped at 045/2134Z (ROP NS-3). The SSR EDAC Counter was cleared seven times (ROP SR-1). The Gyro #3 drift rate has shown no improvement. The rate of change is 36.36 arcseconds/hour/day. The trend still exceeds the desired limit of 10.8 arcseconds/hour/day prior to Gyro #3 reinstallation into the control loop. Gyro #1 Saturation Safing Flag was out of limits (OOL) for one sample (12 Seconds) at the end of a slew at 045/1033Z. HSTAR 6996 was written See Item 1.20 (Proposal 7505). There was a Guide Star Acquisition which resulted in a Fine Lock Backup at 045/1423Z due to a FGS #1 SSLE. The subsequent reacquisiton (1,3,1) resulted in Fine Lock Backup, Fine Lock was achieved using the same primary FGS #1 (1,0,1). HSTART 6997 was written. (See ites 1.5 and 1.9 (Proposals 7475 and 8091)). 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH